Theorem 1. Let Φ = {α}∪Φ^{+} be a data stucture with a bi-variate statistics

Φ → ℕ × X ; p ↦ s(p) = (n, k)

(X is set of labels) such that

s(Φ^{+}) ⊂ ℕ^{+} × X

Set Φ_{n} = {p ∈ Φ|pr_{1}(s(p)) = n} and l(p) = pr_{2}(s(p)) “label of p”. We suppose that there exists a function “return to the father”

d : Φ_{n+1} → Φ_{n}

such that p ↦ (d(p), l(p)) is into. Then, for p ∈ Φ_{n}, the code

(l(p), l(d(p)), l(d^{2}(p)), ⋯ l(d^{n−1}(p)))

is into.